OpenStudy (anonymous):

what is the slope intercept form of the function containing (-1, 5) and slope of -4? y= x+

OpenStudy (anonymous):

would this be y=5x+1

OpenStudy (anonymous):

y=-1 x + 5

OpenStudy (anonymous):

y2-y1=m(x2-x1) y2-(5)=(-4)(x2-(-1)) y2-5=-4(x2+1) y2-5=-4x-4 y2=-4x-4+5 adding +5 for both side y=-4x+1

OpenStudy (shamim):

@AJ01 u did wrong!!!!!!!!!

OpenStudy (shamim):

Required equation of a line is y=mx+b

OpenStudy (shamim):

Given m=-4 x=-1 y=5

OpenStudy (shamim):

Use my equation b=?

OpenStudy (shamim):

Response plz!!!!

OpenStudy (anonymous):

Thanks, so which one is correct?